OpenStudy (anonymous):

The weights of 1250 pumpkins are normally distributed with mean of 28 pounds and a standard deviation of 6 pounds. About how many of the pumpkins weigh more than 34 pounds. Equation will help. 1050 1000 250 200

OpenStudy (kropot72):

This can be solved by using a standard normal distribution table. The z-score for 34 pounds is 1, the reason being that 34 is one standard deviation above the mean of 28 pounds. Can use the table to find the cumulative probability for z = 1.00 and post the result? If you do this we can do the next simple steps.

OpenStudy (kropot72):

You can use the table here if you need it: http://lilt.ilstu.edu/dasacke/eco148/ztable.htm

OpenStudy (kropot72):

@Cierra104 Are you there?

OpenStudy (anonymous):

yes -2.0

OpenStudy (anonymous):

do you subtract than divide or subtract then multiply

OpenStudy (kropot72):

Did you use a standard normal distribution table? Your result is not a probability.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This how i think i need to do it : 1250/28 Find deviation of 6 to 34?

OpenStudy (anonymous):

I don't understand without an equation thats all i need along with equation for finder's deviation. I can figure the rest out.

OpenStudy (kropot72):

Let p be the cumulative probability for z = 1.00. Then the approximate number of pumpkins weighing more than 34 pounds is given by: \[(1-p) \times1250=you\ can\ calculate\ (when\ you\ find\ the\ value\ of\ p)\]

OpenStudy (kropot72):

The z-score is found as follows: \[z=\frac{X-\mu}{\sigma}=\frac{34-28}{6}=1.00\]

OpenStudy (anonymous):

1-2 x 1250=-1250 of 34? or 28

OpenStudy (anonymous):

i got c

OpenStudy (kropot72):

Where did you get the value of -2 from?

OpenStudy (anonymous):

table

OpenStudy (kropot72):

The value of p must be found from the table here, using z = 1.00 http://lilt.ilstu.edu/dasacke/eco148/ztable.htm

OpenStudy (anonymous):

.8341

OpenStudy (kropot72):

Good work! You are correct. The approximate number of pumpkins weighing more than 34 pounds is given by: \[(1.0000-0.8413) \times1250=you\ can\ calculate\]

OpenStudy (anonymous):

i got 198.375 though do i need to divide?

OpenStudy (kropot72):

You are correct. Just round up to 200 and find the correct answer choice is D.

OpenStudy (anonymous):

or round up. do i always use that chart when it comes to deviation equation or does it depend. my teacher doesn't provide me with this at all.

OpenStudy (kropot72):

There are other methods to solve this question. The empirical rule for a normal distribution could be used, for example.

OpenStudy (anonymous):

okay thanks add me as friend will you be my tutor please i have sats coming up soon

OpenStudy (kropot72):

You're welcome :) I am happy to try to help you in future if you post more questions.

OpenStudy (anonymous):

ok great.

OpenStudy (anonymous):

you good @ chem?